OpenStudy (anonymous):

Solve: 16x2 − 81 = 0. x = 4, 9 x = −4, −9 x = 9 over 4, x = −9 over 4 x = 81 over 16, x = −81 over 16

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

note that 16x^2 and 81 are 2 perfect squares so what are the factors of 16x^2 - 81?

OpenStudy (welshfella):

this is the difference of 2 squares a^2 - b^2 = (a + b)(a - b)

OpenStudy (anonymous):

I do believe the answer is D.

OpenStudy (welshfella):

try to factor the expression compare 16x^2 - 81 with a^2 - b^2 and plug the values in

OpenStudy (anonymous):

16^12-81^2?

OpenStudy (welshfella):

what are the square roots of 16x^2 and 81?

OpenStudy (welshfella):

no a and b are the square roots of 16x^2 and 81

OpenStudy (anonymous):

I think the answer is D. Am i correct?

OpenStudy (welshfella):

no

OpenStudy (welshfella):

what number multiplied by itself = 81?

OpenStudy (anonymous):

The answer I belive it is @welshfella make sure I am correct on this lol. B.

OpenStudy (anonymous):

@welshfella is she correct?

OpenStudy (welshfella):

no its not d the important thing is you learn how to solve this

OpenStudy (anonymous):

ok

OpenStudy (welshfella):

ok lets do it another way 16x^2 - 81 = 0 add 81 to both sides 16x^2 = 81 now divide both sides by 16 what do you get?

OpenStudy (anonymous):

I am confused.

OpenStudy (welshfella):